OpenStudy (anonymous):

How can 1/5x − 2 = 1/3x + 8 be set up as a system of equations?

OpenStudy (anonymous):

Are there any answer choices?

OpenStudy (anonymous):

5y − 5x = −10 3y − 3x = 24 5y − 5x = −10 3y + 3x = 24 5y + x = −10 3y + x = 24 5y − x = −10 3y − x = 24

OpenStudy (anonymous):

I think it might be the second one. Let me see if I can find someone to make sure

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

@adrynicoleb @tester97